# Entropy-based random models for hypergraphs

### v2

@@ Abstract
$\langle\sigma_{i}\rangle$ $=\sum_{\begin{subarray}{c}j=1\\ (j\neq i)\end{subarray}}^{N}\langle w_{ij}\rangle=\sum_{\alpha=1}^{L}p_{i\alpha}[\langle h_{\alpha}\rangle-p_{i\alpha}].$ (39)
As further confirmed by Fig. 8 in Appendix A, the approximation provided by Eq. (34) allows us to pose $\langle\sigma_{i}\rangle_{\text{HCM}}\simeq k_{i}^{*}\sum_{\alpha=1}^{L}(h_{\alpha}^{*})^{2}/T^{*}$, $\forall\>i$. Interestingly, as Fig. 5a shows, the HCM overestimates the extent to which any two nodes of the email-Enron data set overlap: in words, such a real-world hypergraph is more compartmentalised than expected.
@@
$\langle k_{i}^{nn}\rangle$ $\simeq\sum_{\alpha=1}^{L}\frac{p_{i\alpha}[\langle h_{\alpha}\rangle+1-p_{i\alpha}]}{\langle k_{i}\rangle}$
$=\frac{\langle\sigma_{i}\rangle+\langle k_{i}\rangle}{\langle k_{i}\rangle}=\frac{\langle\sigma_{i}\rangle}{\langle k_{i}\rangle}+1\simeq\frac{\langle\sigma_{i}\rangle}{\langle k_{i}\rangle}.$ (41)
IV.4 D. Disparity ratio and degree in the projection
@@
$Y_{i}=\sum_{\begin{subarray}{c}j=1\\ (j\neq i)\end{subarray}}^{N}f_{ij}^{2}=\sum_{\begin{subarray}{c}j=1\\ (j\neq i)\end{subarray}}^{N}\frac{w_{ij}^{2}}{\sigma_{i}^{2}},$ (42)
known as disparity ratio and quantifying the (un)evenness of the distribution of the weights constituting the strength of node $i$ over the $\kappa_{i}=\sum_{\begin{subarray}{c}j=1\\ (j\neq i)\end{subarray}}^{N}\Theta[w_{ij}]\equiv\sum_{\begin{subarray}{c}j=1\\ (j\neq i)\end{subarray}}^{N}a_{ij}$ links characterising its connectivity - since $a_{ij}=1$ if nodes $i$ and $j$ share, at least, one hyperedge, $\kappa_{i}$ is the degree of node $i$ in the projection of the hypergraph (see also Fig. 4 and Fig. 5b). Since, under the RHM, $w_{ij}\sim\text{Bin}(L,p^{2})$, we find that $\langle a_{ij}\rangle=1-(1-p^{2})^{L}$, i.e. the expected value of $a_{ij}$ coincides with the probability of observing a non-zero overlap. Under the HCM, instead, $w_{ij}\sim\text{PoissBin}(L,\{p_{i\alpha}p_{j\alpha}\}_{\alpha=1}^{L})$, hence
@@
$\langle Y_{i}\rangle\simeq\sum_{\begin{subarray}{c}j=1\\ (j\neq i)\end{subarray}}^{N}\frac{\langle w_{ij}^{2}\rangle}{\langle\sigma_{i}^{2}\rangle}.$ (45)
Figure 5: Scatter plots between the empirical values $\{\sigma_{i}\}$, $\{\kappa_{i}\}$, $\{Y_{i}\}$, $\{\text{CEC}_{i}\}$ and the expected values $\{\langle\sigma_{i}\rangle_{\text{HCM}}\}$, $\{\langle\kappa_{i}\rangle_{\text{HCM}}\}$, $\langle\{Y_{i}\rangle_{\text{HCM}}\}$, $\{\langle\text{CEC}_{i}\rangle_{\text{HCM}}\}$. The HCM overestimates the extent to which any, two nodes overlap, as well as the CEC; the disparity ratio, instead, is underestimated by it. These results can be understood by considering that the HCM just constrains the degree sequences, hence inducing an ensemble where connections are ‘distributed’ more evenly than observed.
@@
Contrarily to what has been previously observed, the expected value of the disparity ratio cannot be always safely decomposed as a ratio of expected values - not even if the ‘full’ HCM is employed. In fact, while this approximation works relatively well for the contact-primary-school data set, it does not for the email-Enron and the NDC-classes ones (see also Fig. 10 in Appendix C). For this reason, the expected value of the disparity ratio has been evaluated by explicitly sampling the ensemble of incidence matrices induced by the ‘full’ HCM. In any case, as Fig. 5c shows, such a null model underestimates the disparity ratio characterising each node of the email-Enron data set: in words, the empirical overlap between any two nodes is (much) less evenly ‘distributed’ than expected. A similar conclusion can be drawn by considering $\langle\kappa_{i}\rangle=\sum_{i=1}^{N}[1-\prod_{\alpha=1}^{L}(1-p_{i\alpha}p_{j\alpha})]$: as Fig. 5b shows, the degree of the nodes in the projection tend to be significantly smaller than expected, meaning that hyperedges concentrate on fewer edges than expected. This observation is in line with recent works showing the encapsulation and ‘simpliciality’ of real-world hypergraphs LaRock and Lambiotte (2023); Landry et al. (2024).
IV.5 F. Eigenvector centrality
@@
Centrality measures for hypergraphs have been defined as well. An example is provided by the clique motif eigenvector centrality (CEC), defined in Benson (2019) (see also Appendix D): $\text{CEC}_{i}$ corresponds to the $i$-th entry of the Perron-Frobenius eigenvector of $\mathbf{W}$. As Fig. 5d shows, the HCM underestimates the CEC as well: such a result can be understood by considering that the HCM constrains only the degree sequences, hence inducing an ensemble where connections are ‘distributed’ more evenly than observed, an evidence letting the nodes overlap more, thus causing the entries of $\langle\mathbf{W}\rangle$ to be overall larger and less dissimilar, as well as those of its Perron-Frobenius eigenvector.
Figure 6: Top panels: projections of the contact-primary-school, email-Enron and NDC-classes data sets onto the layer of nodes. Bottom panels: validated counterparts of the aforementioned projections: any two nodes are linked if they share a significantly large number of hyperedges. Communities have been detected by running the Louvain algorithm.
IV.6 G. Confusion matrix
@@
Let us, now, consider the set of properties composing the so-called confusion matrix (see also Appendix E). They are known as the true positive rate (TPR), i.e. the percentage of 1s correctly recovered by a given method and whose expected value reads
@@
the specificity (SPC), i.e. the percentage of 0s correctly recovered by a given method and whose expected value reads
@@
the positive predictive value (PPV), i.e. the percentage of 1s correctly recovered by a given method with respect to the total number of 1s predicted by it and whose expected value reads
@@
and the accuracy (ACC), measuring the overall performance of a given method in correctly placing both 1s and 0s and whose expected value reads
@@
Results on the confusion matrix of a number of real-world hypergraphs reveal that the large sparsity of the latter ones makes it difficult to reproduce the TPR and the PPV (see also Table 3 in Appendix E); on the other hand, the capability of the HCM (both in its ‘full’ and approximated version) to reproduce the density of 1s - and, as a consequence, the density of 0s - ensures the SPC to be recovered quite precisely, in turn ensuring the overall accuracy of the model to be large (for an overall evaluation of the performance of the HCM in reproducing real-world hypergraphs, see also Table 4 in Appendix F).
IV.7 H. Community detection
Communities are commonly understood as densely connected groups of nodes. Representing an hypergraph via its incidence matrix allows this statement to be made more precise from a statistical perspective: in fact, the null models discussed so far can be employed to test if any two nodes share a significantly large number of hyperedges - hence can be clustered together, should this be the case.
To this aim, we can adapt the recipe proposed in Saracco et al. (2017) to project bipartite networks and summarised in the following. One, first, computes
@@
Figure 6 shows the result returned by the Louvain algorithm, run on the validated projections: as noticed elsewhere Saracco et al. (2017); Pratelli et al. (2024); Guarino et al. (2024), the detection of mesoscale structures is enhanced if carried out on filtered topologies.
V V. Conclusions
Our paper contributes to current research on hypergraphs by extending the constrained entropy-maximisation framework to incidence matrices, i.e. their simplest, tabular representation. Differently from the currently-available techniques Chodrow (2020), our methodology has the advantage of being analytically tractable, scalable and versatile enough to be straightforwardly extensible to directed and/or weighted hypergraphs.
Beside leading to results whose relevance is mostly theoretical (i.e. the individuation of different regimes for higher-order structures and the estimation of the actual impact of empty and parallel hyperedges on the analysis of empirical systems), our models prove to be particularly useful when employed as benchmarks for real-world systems, i.e. for detecting patterns that are not imputable to purely random effects. Specifically, our results suggest that real-world hypergraphs are characterised by a degree of self-organisation that is absolutely non-trivial (see also Appendix F).
This is even more surprising when considering that our results are obtained under a benchmark such as the HCM, i.e. a null model constraining both the degree and the hyperdegree sequences: since it overestimates the extent to which any two nodes overlap - a result whose relevance becomes evident as soon as one considers the effects that higher-order structures have on spreading and cooperation processes de Arruda et al. (2021); Iacopini et al. (2022); St-Onge et al. (2022) - our future efforts will be directed towards the analysis of benchmarks constraining non-linear quantities such as the co-occurrences between nodes and/or hyperedges.
VI VI. Acknowledgements

### v3

@@ Results
$\langle\sigma_{i}\rangle$ $=\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{N}\langle w_{ij}\rangle=\sum_{\alpha=1}^{L}p_{i\alpha}[\langle h_{\alpha}\rangle-p_{i\alpha}].$ (63)
As further confirmed by Fig. 11 in the Appendix, the approximation provided by Eq. (58) allows us to pose $\langle\sigma_{i}\rangle_{\text{HCM}}\simeq k_{i}^{*}\sum_{\alpha=1}^{L}(h_{\alpha}^{*})^{2}/T^{*}$, $\forall\>i$. Interestingly, as Fig. 9a shows, the HCM overestimates the extent to which any two nodes of the email-Enron data set overlap: in words, such a real-world hypergraph is more compartmentalised than expected.
@@
$\langle k_{i}^{nn}\rangle$ $\simeq\sum_{\alpha=1}^{L}\frac{p_{i\alpha}[\langle h_{\alpha}\rangle+1-p_{i\alpha}]}{\langle k_{i}\rangle}=\frac{\langle\sigma_{i}\rangle+\langle k_{i}\rangle}{\langle k_{i}\rangle}=\frac{\langle\sigma_{i}\rangle}{\langle k_{i}\rangle}+1\simeq\frac{\langle\sigma_{i}\rangle}{\langle k_{i}\rangle}.$ (65)
Disparity ratio and degree in the projection
@@
$Y_{i}=\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{N}f_{ij}^{2}=\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{N}\frac{w_{ij}^{2}}{\sigma_{i}^{2}},$ (66)
known as disparity ratio and quantifying the (un)evenness of the distribution of the weights constituting the strength of node $i$ over the $\kappa_{i}=\sum_{\begin{subarray}{c}j=1\\ (j\neq i)\end{subarray}}^{N}\Theta[w_{ij}]\equiv\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{N}a_{ij}$ links characterising its connectivity - since $a_{ij}=1$ if nodes $i$ and $j$ share, at least, one hyperedge, $\kappa_{i}$ is the degree of node $i$ in the projection of the hypergraph (see also Fig. 8 and Fig. 9b). Since, under the RHM, $w_{ij}\sim\text{Bin}(L,p^{2})$, we find that $\langle a_{ij}\rangle=1-(1-p^{2})^{L}$, i.e. the expected value of $a_{ij}$ coincides with the probability of observing a non-zero overlap. Under the HCM, instead, $w_{ij}\sim\text{PoissBin}(L,\{p_{i\alpha}p_{j\alpha}\}_{\alpha=1}^{L})$, hence
@@
$\langle Y_{i}\rangle\simeq\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{N}\frac{\langle w_{ij}^{2}\rangle}{\langle\sigma_{i}^{2}\rangle}.$ (69)
@@
Contrarily to what has been previously observed, the expected value of the disparity ratio cannot be always safely decomposed as a ratio of expected values - not even if the ‘full’ HCM is employed. In fact, while this approximation works relatively well for the contact-primary-school data set, it does not for the email-Enron and the NDC-classes ones (see also Fig. 13 in the Appendix). For this reason, the expected value of the disparity ratio has been evaluated by explicitly sampling the ensemble of incidence matrices induced by the ‘full’ HCM. In any case, as Fig. 9c shows, such a null model underestimates the disparity ratio characterising each node of the email-Enron data set: in words, the empirical overlap between any two nodes is (much) less evenly ‘distributed’ than expected. A similar conclusion can be drawn by considering $\langle\kappa_{i}\rangle=\sum_{i=1}^{N}\left[1-\prod_{\alpha=1}^{L}(1-p_{i\alpha}p_{j\alpha})\right]$: as Fig. 9b shows, the degree of the nodes in the projection tend to be significantly smaller than expected, meaning that hyperedges concentrate on fewer edges than expected. This observation is in line with recent works showing the encapsulation and ‘simpliciality’ of real-world hypergraphs [42, 43].
Figure 9: Scatter plots between the empirical and the expected values for the email-Enron dataset. More in detail, $\{\sigma_{i}\}$ vs. $\{\langle\sigma_{i}\rangle_{\text{HCM}}\}$ (panel a), $\{\kappa_{i}\}$ vs. $\{\langle\kappa_{i}\rangle_{\text{HCM}}\}$ (panel b), $\{Y_{i}\}$ vs. $\langle\{Y_{i}\rangle_{\text{HCM}}\}$ (panel c), and $\{\text{CEC}_{i}\}$ vs. $\{\langle\text{CEC}_{i}\rangle_{\text{HCM}}\}$ (panel d). The HCM overestimates the extent to which any, two nodes overlap, as well as the CEC; the disparity ratio, instead, is underestimated by it. These results can be understood by considering that the HCM just constrains the degree sequences, hence inducing an ensemble where connections are ‘distributed’ more evenly than observed.
Eigenvector centrality
@@
Centrality measures for hypergraphs have been defined as well. An example is provided by the clique motif eigenvector centrality (CEC), defined in [44] (see also Section C of the Appendix): $\text{CEC}_{i}$ corresponds to the $i$-th entry of the Perron-Frobenius eigenvector of $\mathbf{W}$. As Fig. 9d shows, the HCM underestimates the CEC as well: such a result can be understood by considering that the HCM constrains only the degree sequences, hence inducing an ensemble where connections are ‘distributed’ more evenly than observed, an evidence letting the nodes overlap more, thus causing the entries of $\langle\mathbf{W}\rangle$ to be overall larger and less dissimilar, as well as those of its Perron-Frobenius eigenvector.
Confusion matrix
@@
Let us, now, consider the set of indices constituting the so-called confusion matrix (see also Section D of the Appendix). They are intended to quantify the capability of a given network model in reproducing microscopic properties such as the position of $1$s and 00s by explicitly comparing their empirical location with the one expected under the chosen model. They are named true positive rate (TPR), i.e. the percentage of $1$s correctly recovered by a given method, whose expected value reads
@@
specificity (SPC), i.e. the percentage of 0s correctly recovered by a given method, whose expected value reads
@@
positive predictive value (PPV), i.e. the percentage of $1$s correctly recovered by a given method with respect to the total number of $1$s predicted by it, whose expected value reads
@@
accuracy (ACC), measuring the overall performance of a given method in correctly placing both $1$s and 00s, whose expected value reads
@@
Results on the confusion matrix of a number of real-world hypergraphs reveal that the large sparsity of the latter ones makes it difficult to reproduce the TPR and the PPV (see also Table 2 in the Appendix); on the other hand, the capability of the HCM (both in its ‘full’ and approximated version) to reproduce the density of $1$s - and, as a consequence, the density of 00s - ensures the SPC to be recovered quite precisely, in turn ensuring the overall accuracy of the model to be large (for an overall evaluation of the performance of the HCM in reproducing real-world hypergraphs, see also Table 3 in Section E of the Appendix).
Figure 10: Validated vs. non-validated ’hypergraph to network’ projections of empirical datasets. Panels a, b and c: projections of the contact-primary-school, email-Enron and NDC-classes data sets onto the layer of nodes. Panels d, e and f: validated counterparts of the aforementioned projections: any two nodes are linked if they share a significantly large number of hyperedges. Communities have been detected by running the Louvain algorithm.
Community detection
Communities are commonly understood as densely connected groups of nodes. Representing an hypergraph via its incidence matrix allows this statement to be made more precise from a statistical perspective: in fact, the null models discussed so far can be employed to test if any two nodes share a significantly large number of hyperedges - hence can be clustered together, should this be the case. In other words, it is possible to devise a ‘validation procedure’ that filters the projection described by the matrix $\mathbf{W}$ by removing the entries that do not satisfy the requirement above.
To this aim, we can adapt the recipe proposed in [24] to project bipartite networks and summarised in the following. One, first, computes
@@
Fig. 10 shows the partitions returned by the Louvain algorithm run on the validated projections: as noticed elsewhere [24, 46, 47], the detection of mesoscale structures is enhanced if carried out on filtered topologies.
Conclusions
Our paper contributes to current research on hypergraphs by extending the constrained entropy-maximisation framework to incidence matrices, i.e. their simplest, tabular representation. Differently from the currently-available techniques [9], our methodology has the advantage of being analytically tractable, scalable and versatile enough to be straightforwardly extensible to directed and/or weighted hypergraphs.
Beside leading to results whose relevance is mostly theoretical (i.e. the individuation of different regimes for higher-order structures and the estimation of the actual impact of empty and parallel hyperedges on the analysis of empirical systems), our models prove to be particularly useful when employed as benchmarks for real-world systems, i.e. for detecting patterns that are not imputable to purely random effects. Specifically, our results suggest that real-world hypergraphs are characterised by a degree of self-organisation that is absolutely non-trivial (see also Section E of the Appendix).
This is even more surprising when considering that our results are obtained under a benchmark such as the HCM, i.e. a null model constraining both the degree and the hyperdegree sequences: since it overestimates the extent to which any two nodes overlap - a result whose relevance becomes evident as soon as one considers the effects that higher-order structures have on spreading and cooperation processes [48, 49, 50] - our future efforts will be directed towards the analysis of benchmarks constraining non-linear quantities such as the co-occurrences between nodes and/or hyperedges.
[…]
Acknowledgements